OpenStudy (anonymous):

A cooler contains 7 cans of soda. 4 colas, 2 orange, and 1 cherry. Two cans are selected at random without replacement. Find the probability that at least one is cherry. Please help me with the steps.

OpenStudy (anonymous):

is it (1- 13/14)= 1/14

OpenStudy (anonymous):

No I tried that. Apparently there are a lot more steps than I though :(

OpenStudy (anonymous):

i hav used 1 - (c(13,2)/c(14,2))

OpenStudy (anonymous):

The answer was 2/7 but I can not figure out why. I have a test coming up and this is the only part I had problems with

OpenStudy (anonymous):

as per my calc it may be 1/7 (i worngly calculated to 1/14)

OpenStudy (anonymous):

I got it now. but will write it out for future reference. Let P(A) be Exactly one soda is cherry. Let P(B) be Both are Cherry. Break P(A) down into P(C) The first is cherry, second is not. P(D) The first is not cherry and the second is.

OpenStudy (anonymous):

To obtain P(C) Determine how many can be Cherry which is 1/7 and if second is picked how many are not Cherry which is 6/6. 1/7 X 6/6= 6/42 So P(C) and P(D) are 6/42

OpenStudy (anonymous):

P(A) = 6/42 + 6/42 = 12/42 P(B) = 0 because if you pick a cherry the first time, you can not pick another the second. So the answer is 12/42 or 2/7

OpenStudy (anonymous):

Thank you for the help though